if X~B(n₁ = 5, p = 1/4) and Y~B(n = 8, p = 1/4).Then what is the distribution of X+Y?


1
Expert's answer
2022-06-14T00:14:00-0400

If "X \sim B(n, p)" and "Y \sim B(m, p)" are independent binomial variables with the same probability "p," then "X + Y"  is again a binomial variable; its distribution is "Z=X+Y \sim B(n+m, p)."

Given "n_1=5, n=8, p=1\/4."

Then "Z=X+Y \sim B(n_1+n=13, p=1/4)."